casey receives $ 336 per year in simple interest from 3 investments. part is invester at 8%, part 9%, and part at 10%. there is $500 more invested at 9% than at 8%. The amount invested at 10% is 3 times the amount invested at 9%. Fine the amount invested at each rate.
8%________
9%________
10%_______

Answer by stanbon(48568)

casey receives $ 336 per year in simple interest from 3 investments. part is invester at 8%, part 9%, and part at 10%. there is $500 more invested at 9% than at 8%. The amount invested at 10% is 3 times the amount invested at 9%. Fine the amount invested at each rate.
8%________
9%________
10%_______
------------------
Let amt. invested at 8% be "x".
Then amt. invested at 9% is "x+500"
And amt. invested at 10% is "3x+1500"
-------------
EQUATION:
0.08x + 0.09(x+500) + 0.10(3x+1500) = 336
Solve for x
0.08x + 0.09x + 45 + + 0.30x +150 = 336
0.47x = 141
x = $300 (amount invested at 8%)
x+500 = $800.00 (amount invested at 9%
3x+1500 = $2400.00 (amount invested at 10%)
================================================
Cheers,
Stan H.
===============